During the past period, 65,000 units of hinges were made and the materialsquantity variance was $50,000, unfavorable. The standards for direct materials in making this hinge are 40 ounces, per unit, at 0.50 per ounce. What is the number of ounces of direct material used during this period?
Answer #1

Actual ounces used = 2700000 ounces

Material Quantity Variance

(

Standard Quantity

-

Actual Quantity

)

x

Standard Rate

(

2600000

-

2700000

)

x

$       0.50

-50000

Variance

$       50,000.00

Unfavourable-U

MQV= (SQ-AQ) x SR

-50000=(2600000-AQ) x 0.50

-50000= (1300000-0.50AQ)

(1300000+50000)=0.50AQ

AQ=1350000/0.50

AQ=2700000

.

Where AQ= Actual Quantity

SR= Standard rate

SQ=Standard quantity

And

MQV = material quantity variance